OpenStudy (anonymous):

Guy riding a bike accelerates for 8 seconds. He covers 60 meters. Using the formula d=vt+1/2at^2 calculate his acceleration.

OpenStudy (anonymous):

Alright so basically plugging in the numbers you'd think right

OpenStudy (anonymous):

Look 60m=1/2(a)64s^2

OpenStudy (anonymous):

Units don't cancel out

OpenStudy (anonymous):

60m=32s^2a

OpenStudy (anonymous):

I mean wtf

OpenStudy (anonymous):

a=(m/s^2)

OpenStudy (anonymous):

This is such a paradox

OpenStudy (anonymous):

No way to derive units for acceleration with the given kinematics equation which holds true

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

Hey dude what do you think

OpenStudy (anonymous):

lol you stuck with this paradox too

OpenStudy (anonymous):

don't worry you are not the only one

OpenStudy (anonymous):

The only way to do this is to assume "a" doesn't bear any value.

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Merely a means to represent imaginary number which follows from the function of this equation

OpenStudy (anonymous):

\[v=m/s \]

OpenStudy (anonymous):

But from the principles of physics which assigns unit to every different letter this is a paradox

OpenStudy (anonymous):

eh?

OpenStudy (anonymous):

\[d=m\] \[a=m/s^2\]

OpenStudy (anonymous):

Then the prospect of using a is misconceived because a in this equation doesn't assume the units m/s^2 No

OpenStudy (anonymous):

Instead I am left with the presupposition that a will have the units and must

OpenStudy (anonymous):

60m/32s^2=a

OpenStudy (anonymous):

total distance traveled=initial distance+(velocity)(time)+0.50(acceleration)(time)^2

OpenStudy (anonymous):

1.875m/s^2

OpenStudy (anonymous):

In this case the initial distance and initial velocity is ignored as the cyclist starts from the state of being at rest

OpenStudy (anonymous):

\[m=m+(\frac{ m }{ s }*s)+0.50*\frac{ m }{ s^2 }*s^2\]

OpenStudy (anonymous):

total m=m+m+0.50*m

OpenStudy (anonymous):

you are wrong

OpenStudy (anonymous):

you can't assume a to have (m/s^2) internally.

OpenStudy (anonymous):

the units of a are assigned subsequently from the calculation. Assigning such units without solid foundation only serves to confuse the user.

OpenStudy (anonymous):

acceleration is m/s^2

OpenStudy (anonymous):

Look 60m=32s^2a a=1.875m/s^2

OpenStudy (anonymous):

This is correct.

OpenStudy (anonymous):

But if you assume a=(s/m^2) that gets you nowhere.

OpenStudy (anonymous):

Try it for yourself with units to be accorded

OpenStudy (anonymous):

This is where the teaching of physics fails.

OpenStudy (anonymous):

Each representative is not to be assumed to have the units but rather guidelines for what they stand for and what the user should expect to get following the calculations.

OpenStudy (anonymous):

acceleration is velocity over time. Acceleration is how fast velocity is changing with time since velocity/time is acceleration. and velocity is m/s and time is s (m/s)*(1/s)=m/s^2

OpenStudy (anonymous):

By no means assign the units prior to calculation as it throws of the equation.

OpenStudy (anonymous):

But it's interesting.

OpenStudy (anonymous):

so we know acceleration is m/s^2 because acceleration is known from how fast velocity is changing. So they plug a into your equation and the units cancels.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

a=v/t came first before d=d1+vt+1/2)at^2

OpenStudy (anonymous):

When a is imaginary you can't even assign m/s^2

OpenStudy (anonymous):

However when a is known you can naturally assign the units however "naturally" as in being circumstantial and not at the hands of calculator

OpenStudy (anonymous):

there is no imaginary. there is an estimated A.

OpenStudy (anonymous):

hypothetical A

OpenStudy (anonymous):

so that hypothetical a doesn't assume units in itself

OpenStudy (anonymous):

You know what I mean? So many amateurs like me are presupposing upon seeing the hypothetical a or v or even d that they have units in themselves but this assumption is totally off

OpenStudy (anonymous):

Yes, it can if velocity and time can too?

OpenStudy (anonymous):

and acceleration is derived from them

OpenStudy (anonymous):

This is quite contradictory from what physics teachers force you to assume.

OpenStudy (anonymous):

we use seconds for time and units for speed why not use acceleration by dividing the units out

OpenStudy (anonymous):

Nothing imaginary about that. It is not abstract like math.

OpenStudy (anonymous):

I think my best bet is whenever you see a letter that makes a reference to set of units you need to assume that they in themselves do not possess anything of the sort however it's a guideline instructing the user of that equation to assume that the final work will be concordant with the standard form.

OpenStudy (anonymous):

Yeah physics is very paradoxical sometimes but that is our passion right? XD

OpenStudy (anonymous):

Yes, but it is also inaccurate in a lot of ways.

OpenStudy (anonymous):

PV=nRT for ideal gas law n does not make sense for me also.

OpenStudy (anonymous):

that is my opinion on some part of physics, but I think the acceleration actually does make sense.

OpenStudy (anonymous):

Yeah. Sometimes physics ignores the principles of math and heavily relies on intuitive reasoning which often fails. But this is the reality.

OpenStudy (anonymous):

The prospect of entirely relying on single solution upon a problem of kind shows it all. Unfortunate but yet convenient.

OpenStudy (anonymous):

Yep, that is the difference between experimental and theoretical physicists. They try to prove each other wrong. Well I got to sleep now. See you another day.

OpenStudy (anonymous):

Yeah Interesting talk my friend;) Good to see you.